OpenStudy (anonymous):

How do I solve a differential equation if the equation becomes undefined when I plug in the initial value?

OpenStudy (anonymous):

heyllo joseph hru?

OpenStudy (anonymous):

Show the equation :D If it's undefined at the initial value, then you probably integrated incorrectly :D

OpenStudy (anonymous):

dx/dy = (y-1)^2

OpenStudy (anonymous):

x = 1/3(y-1)^3 + C What are the initial conditions?

OpenStudy (anonymous):

y(0) = 1

OpenStudy (anonymous):

C = 0

OpenStudy (anonymous):

x = 1/3 (y-1)^3

OpenStudy (anonymous):

can you show your work in getting the solution to the differential equation?

OpenStudy (anonymous):

y(0) = 1 x = 0 y = 1 0 = 1/3(1-1)^3 +C 0 = 0 +C C = 0

OpenStudy (anonymous):

excuse me! dy/dx =...

OpenStudy (anonymous):

Oh:\[dx/dy = (y-1)^2\] \[\int\limits_{0}^{x}x dx = \int\limits_{0}^{y} (y-1)^2 dy\] Take the antiderivative (in this case, it works) :D \[x = (y-1)^3/3\]

OpenStudy (anonymous):

+C

OpenStudy (anonymous):

DON'T FORGET THE +C

OpenStudy (anonymous):

k

OpenStudy (anonymous):

excuse me, i made a mistake, the dy and dx are flipped from what I wrote. Is it OK?

OpenStudy (anonymous):

No, then the equation is completely different: \[dy/dx = (y-1)^2\] Limits using initial conditions \[\int\limits_{1}^{y} dy/(y-1)^2 = \int\limits_{0}^{x} xdx\] \[\int\limits_{1}^{y} (y-1)^{-2} dy = x\] Antidifferentiate again:\[-(y-1)^{-1} - - (0)^{-1} = x \] No solution for this diff eq.